How many grams of AgBr will form when 45.0 mL of 0.842 M CaBr2 react?

CaBr2(aq) + 2 AgNO3(aq) → 2 AgBr(s) + Ca(NO3)2(aq)
1
Expert's answer
2020-11-23T06:55:00-0500

"45mL = 0.045L"

"n = C*V"

"n(CaBr_2)=0.842*0.045mol"

"n(AgBr)= 2*n(CaBr_2)="

"=2*0.842*0.045mol=0.07578mol"

"m=n*M"

"M(AgBr)= 188g\/mol"

"m(AgBr)= 188*0.07578g=14.24664g"


Answer: "m(AgBr)=14.24664g"